Researchers will create a confidence interval for the proportion of all first-grade children who are overweight. The chance that this method will provide a correct result is determined by the

a. sample size.

b. level of confidence

c. sample mean.

d. population standard deviation.
Answer #1

The level of confidence is given by: 100(1-\alpha)

Based on 100(1-\alpha)\% confidence interval, we can say that if confidence interval of the same sample size are formed, then the true proportion will lie in the interval 100(1-\alpha)\% times.

Hence option B.